OpenStudy (anonymous):

the area of a rectangle is 600 in 2. find the dimensions of the rectangle if they are in a ratio of 2:3

OpenStudy (anonymous):

based on quadratics

OpenStudy (anonymous):

what equation do i make besides xy=600?

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

you are given with dimensions ratio..

OpenStudy (anonymous):

So, let the dimensions be : \(2x\) and \(3x\)..

OpenStudy (anonymous):

Always, use the given ratio to define your dimensions..

OpenStudy (anonymous):

So, let the length be : \(3x\) and the width is : \(2x\)..

OpenStudy (anonymous):

Use this and find out value of \(x\)..

OpenStudy (anonymous):

It will confuse you, for length, take \(l\) and for width, take \(w\)..

OpenStudy (anonymous):

3x*2x=600?

OpenStudy (anonymous):

yep, good..

OpenStudy (anonymous):

\[l = 3x \\ w = 2x \\ lw = 600 \\ 3x \cdot 2x = 600 \\ 6x^2 = 600\]

OpenStudy (anonymous):

20in by 30in

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

See, take the ratio of dimensions now, you will surely get 2/3..

OpenStudy (anonymous):

\[\frac{20}{30} = \frac{2}{3}\]

OpenStudy (anonymous):

Do you have options with you?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Like 4 answer choices from which you have to choose any one?

OpenStudy (anonymous):

do 600 divided by 5 since it is 2+3 600/5 would be 120 so that would be x 120(2): 120(3) = 240: 360 but then you would divide 240 and 360 by 2 since it is a rectangle so it would be 120:180

OpenStudy (anonymous):

do you understand??